if 1 joule of work is done to move 1 coloumb from one point to another, what is the potential difference

Grade:10

2 Answers

Piyush Kumar Behera
4 years ago
Potential is defined as the work done per unit charge.
So for the above question work done is  1J and charge is 1 coloumb.
Hence the potential difference is \frac{1J}{1C}= 1Volt
Hence the answer is 1 volt
 
I hope it helps!!
Khimraj
4 years ago

work done (W) = 1J

Charge       (Q) = 1C

Potential difference (V) =?

V = W/Q

V = 1 J ÷ 1 C = 1 V

…........................................................